please help

dy 1. Solve the initial value problem, dr2 8y-16% y(0) = 5, 7(0) = 0. de
Answer #1

trial coln the day ane 24 - sy = 16 e 2x let, Y = Aema be the homogeons So, m-am - 8 = 0 OX, mr- 4m +2m-8=0 ON (m-4) (m+2) =